OpenStudy (emilyjones284):

Solve the equation. 2x= radical 20+2x

OpenStudy (anonymous):

isolate the radical i think then square both sides

OpenStudy (anonymous):

2x= radical 20+2x 0=sqrt 20 It looks like no solution maybe i doing it wrong.

OpenStudy (anonymous):

i dbld checked and there is no solution. It false

OpenStudy (anonymous):

unles you meant 2x= radical (20+2x)

OpenStudy (anonymous):

(2x)^2=20+2x 2x*2x=20+2x and i am looking more at it lol.

OpenStudy (anonymous):

4x^2=20+2x

OpenStudy (anonymous):

you gonna have to use a quadratic equation then (2x)^2=20+2x 4x^2-2x-20=0 get all in the left hand side making the right 0

OpenStudy (anonymous):

here you'd use the quadratic formula.

OpenStudy (emilyjones284):

thanks!